OpenStudy (anonymous):

Solve for x: 4 over 3x + 2 over 3 = 6

OpenStudy (anonymous):

Is it 4/3x+2/3=6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4+2x=18x 4=16x x=1/4